OpenStudy (anonymous):

(-3x10^-5)(2x10^3) evaluate and put in standard form

OpenStudy (anonymous):

\[(-3 \times \frac{ 1 }{ 10^5 })(2 \times 10^3) \] can you solve further?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

\[(\frac{ -3 }{ 10^5 })(2 \times 10^3) \] \[(\frac{ -3 \times 2 \times 10^3 }{ 10^5 })\] \[(\frac{ -6 \ }{ 10^2 })\] \[-6 \times 10^{-2}\]